# Periodic Cosmic String Formation and Dynamics

### v1

@@ 5.2.1 Small Initial $\mathcal{J}_{0}$
We begin by considering the the scalar field dynamics for small $\mathcal{J}_{0}$ on an expanding background cosmology in both two and three spatial dimensions.
@@ 5.2.2 Large Initial $\mathcal{J}_{0}$
The comoving vortex number density begins to decrease when the domain boundaries on which they form, and the counter-rotating regions that sustain the domain boundaries, also begin to disappear. This is most evident from the bottom panel of Fig. 26, where $\mathcal{V}_{\rm coun}/\mathcal{V}_{\rm co}$ also decreases dramatically around $a\approx 10$. Due to the remaining $\langle j_{0}\rangle\simeq\mu v^{2}$ (resulting from the large $\mathcal{J}_{0}$), the system then enters a phase of stable circular motion dominated by large $K_{\theta}$, together with a non-relativistic radiation component $K_{\rho}\gtrsim G_{\rho}$. However, there is no parametric difference between the size of $K_{\theta}$ vs. $K_{\rho}$, $G_{\rho}$, and $G_{\theta}$ at this moment, due to continuous production of radiation through vortex dynamics. Since $K_{\theta}$ would now redshift as $a^{-2D}$, while $K_{\rho}$ would redshift like matter ($a^{-D}$) and $G_{\theta}$ would redshift like radiation ($a^{-(D+1)}$), this late kination-dominated period, if it exists, would be very short. Note also that, even though $K_{\theta}$ is the largest component of the energy density during much of the earlier evolution shown in Fig. 26 during which vortices are present (spanning around a decade of growth in $a$), the scaling $K_{\theta}\propto a^{-8/3}$ (i.e., $a^{-4D/3}$ for $D=2$) is not a period of kination (which would be $a^{-2D}$). [^35] In fact, in $D=3$, $a^{-4D/3}$ would be $a^{-4}$, the same scaling as radiation.
@@ 6 Phenomenological Implications: Periodic Gravitational-Wave Bursts
[…]
@@ 7 Conclusion and Remarks
The results presented in this paper may also be useful for understanding preheating after inflation. [^37] We thank Sergey Sibiryakov for discussions on this point and pointing out Tkachev:1998dc , in which string formation was first observed in the study of preheating after inflation. Separately, an interesting question for future work is whether the dynamics we discovered in this paper survive if the $U(1)$ is weakly gauged, or if the $U(1)$ is extended to a larger symmetry group (e.g., the case of the Standard Model electroweak sector) Dufaux:2010cf ; Rajantie:2000nj ; Amin:2014eta . Would defects form periodically if the Standard Model Higgs were displaced in the radial direction, and would those dynamics remain synchronized across Hubble patches?
@@ Acknowledgments
We thank Will East, Sung-Sik Lee, Sergey Sibiryakov, and Zach Weiner for helpful discussions, and Anson Hook for comments on the draft. This work used anvil at Purdue University through allocation PHY240191 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program access , which is supported by National Science Foundation (NSF) Grants OAC-2138259, -2138286, -2138307, -2137603, and -2138296. J.H. would like to thank KITP for hospitality during the completion of this work. This research was supported in part by NSF Grant PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). M.A.F. thanks the Aspen Center for Physics for hospitality during early stages of this work, supported by NSF Grant PHY-2210452. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities.
@@ Appendix A Concise Review of Vortices in Condensed Matter Systems
[…]
first discussed in 1985PhRvL..55.2887H ; 1993PhyA..200…42T , is an effective description of vortex interactions in the Ginzburg–Landau model in 2D (classical XY model). Here, $\vec{R}_{i}=\left\{X_{i},Y_{i}\right\}$ is the location of the vortex, while the circulation $K=\pm 2\pi/m_{*}$ for vortices ($+$) and anti-vortices (--). It was noted in 1993PhyA..200…42T that the first term in the action is a Berry phase and that this action resembles the action describing the motion of a charged particle (in the XY-plane) in an external magnetic field pointing in the Z-direction (see Tong:2016kpv for a detailed discussion about the connection with the quantum Hall effect). The resulting equation of motion of a vortex $i$ in the background of all the other vortices can be found to be
@@
Particle–vortex duality sheds light on this problem Peskin:1977kp ; Dasgupta:1981zz ; Karch:2016sxi ; Senthil:2018cru ; beekman2011electrodynamics . In two dimensions, particle–vortex duality can be established between the Abelian Higgs model and the XY model. Specifically, the Coulomb phase of the Abelian Higgs model is dual to the Higgsed phase of the XY model. The charged Higgs in the Abelian Higgs model is dual to the vortex in the XY model, and the gauge field is dual to the velocity field through Senthil:2018cru (see footnote 19 for the definition of the Levi-Civita tensor in 2+1-dimensions)
@@ Appendix B Simulation Details
respectively. The gradients of the former, just like $\theta$, are ill-defined at vortex locations, leading to numerical artifacts also in the gradient energy density $g_{\theta}$. The equations are discretized using fourth-order accurate finite-difference stencils in conjunction with fourth-order accurate Runge-Kutta time stepping. The grid spacing is $\Delta x$ and related to the timestep size $\Delta t$ by the Courant factor: $\Delta t=\Delta x/2$. Periodic boundary conditions are imposed in both two- and three-dimensional settings, where the box with volume $(2L)^{D}$ has linear size $L=10\mu^{-1}$ and $L=5\mu^{-1}$ for all two-dimensional simulations on a Minkowski and expanding FLRW background spacetime, respectively. Kreiss–Oliger numerical dissipation is used for numerical stability. This implies that modes of wavelength comparable to the grid spacing, $\lambda\sim\mathcal{O}(1)\Delta x$, are damped efficiently and removed from the evolution. The scale-invariant Gaussian random field, entering the initial conditions described in detail in Sec. 2.4, is constructed using the Box–Muller approach. Lastly, vortices are located using the algorithm introduced in Fleury:2015aca . In two spatial dimensions ($D=2$), this unambiguously identifies the vortices piercing a cell. In three spatial dimensions ($D=3$), we search for strings through only a single face of a given cell (say, the face orthogonal to the $+x$-direction) to find the following lower bound on the string length density $\xi$: this density $\xi=\Delta x/V_{b}(N_{S}/N_{T})$ is obtained by comparing the total number of cells, $N_{T}$, to the total number of cells with a face pierced by a string, $N_{S}$, where the total simulation volume is $V_{b}$. This generally underestimates the length density, since the length of a string’s segment associated with a cell is at least [^39] Because the string cores are well-resolved, the string cannot be highly curved within a single cell. While there may be a small, $\mathcal{O}(1)$ inaccuracy in these per-cell length bounds owing to string curvature, this is likely a negligible effect. $\Delta x$. In principle, the length of this segment could be as long as the simulation box’s linear size $2L$ if the string is aligned with the faces. However, in practice, the string network lacks order on scales comparable to the box size. Therefore, the density $\xi$ is likely at most a factor of a few below the true string length density. In those cases with complete string annihilation, the absence of strings piercing all six faces of each cell in the simulation domain was checked.
@@
Subtleties arise when determining the critical radial displacements $\rho_{c}$ as a function of $\mathcal{J}_{0}$ shown in Fig. 12. Not only could the threshold depend on resolution as outlined in the above discussion, but due to the small number density of vortices for $\rho_{i}\gtrsim\rho_{c}$, the critical threshold $\rho_{c}$ may depend on the noise realization $\mathcal{N}(\bm{x})$. Precisely identifying this threshold requires considering an average over a series of noise realizations, as well as varying $\sigma_{\rho_{i}}$. In this work, we simply obtain approximate thresholds as follows. At fixed $\mathcal{J}_{0}$, we perform a suite of simulations spanning a range of initial displacements $\rho_{i}$ from just below to just above the threshold $\rho_{c}$. For each such displacement, we consider only a single noise realization and fix the variance to $\sigma_{\rho_{i}}=10^{-6}\rho_{i}$. We then estimate the threshold as the mean $\rho_{c}\approx(\rho_{i}^{<}+\rho_{i}^{>})/2$, between the initial displacement $\rho_{i}=\rho_{i}^{<}$, which marginally produced vortex pairs, and $\rho_{i}=\rho^{>}_{i}$, which lead to no vortices throughout its evolution. The errorbars shown in Fig. 12 are then simply ±$\pm 2(\rho_{i}^{>}-\rho_{i}^{<})$. Importantly, we do not strictly rule out that vortices may form for $\rho_{\text{min}}<\rho_{i}<\rho_{i}^{<}$, due to the subtleties listed above.
@@
[…]
@@
To test our cosmological implementation, we solve for the evolution of the zero-mode $\rho_{0}$ governed by (10), augmented by the appropriate Hubble friction in two spatial dimensions, as shown in (44). This simpler set of ordinary differential equations can be readily solved using standard numerical integration techniques. The solution can be compared with the early-time (i.e., before unstable modes acquire a significant amplitude) solution of a cosmological simulation solving the full complex scalar field equations on a radiation-dominated FLRW background spacetime. For the purposes of this comparison, we set $\sigma_{\rho_{i}}=10^{-14}\rho_{i}$ to ensure the unstable modes are sourced at sufficiently small amplitudes. In Fig. 28, we compare these two solutions. Clearly, at early times these two solutions are in excellent agreement. Furthermore, with increasing resolution of our fully nonlinear numerical methods, the difference converges at the expected fourth order. To convince ourselves that the vortex annihilation and transition to the radiation state at late times during the cosmological evolution is not due to insufficient resolution, we show the relevant energy densities in the left panel of Fig. 28. Evidently, the densities are converged in the relevant regime, and the transition to radiation is independent of resolution. The curves only begin deviating from each other once $(a\mu)^{-1}$ approaches the grid scale. In the bottom panel of Fig. 28, we show the change of the vortex density with resolution. The periodicity is a consistent feature across resolutions, while the overall density may vary from burst to burst between resolutions (for reasons outlined above). Lastly, the bursts of vortex production in different-resolution simulations experience a mild dephasing between early and lates times. This originates both from an initial phase-offset at the time of the first burst, as well as numerical drift accumulated throughout the simulations. However, the total dephasing of the bursts accumulates to $\sim\mathcal{O}(1)$ radians across $\sim\mathcal{O}(100)$ bursts.

### v2

@@ 5.2.1 Small Initial $\mathcal{J}_{0}$
We begin by considering the scalar field dynamics for small $\mathcal{J}_{0}$ on an expanding background cosmology in both two and three spatial dimensions.
@@ 5.2.2 Large Initial $\mathcal{J}_{0}$
The comoving vortex number density begins to decrease when the domain boundaries on which they form, and the counter-rotating regions that sustain the domain boundaries, also begin to disappear. This is most evident from the bottom panel of Fig. 26, where $\mathcal{V}_{\rm coun}/\mathcal{V}_{\rm co}$ also decreases dramatically around $a\approx 10$. Due to the remaining $\langle j_{0}\rangle\simeq\mu v^{2}$ (resulting from the large $\mathcal{J}_{0}$), the system then enters a phase of stable circular motion dominated by large $K_{\theta}$, together with a non-relativistic radiation component $K_{\rho}\gtrsim G_{\rho}$. However, there is no parametric difference between the size of $K_{\theta}$ vs. $K_{\rho}$, $G_{\rho}$, and $G_{\theta}$ at this moment, due to continuous production of radiation through vortex dynamics. Since $K_{\theta}$ would now redshift as $a^{-2D}$, while $K_{\rho}$ would redshift like matter ($a^{-D}$) and $G_{\theta}$ would redshift like radiation ($a^{-(D+1)}$), any late period in which the dominant energy-density component of the scalar field scales as it would in a kination-dominated epoch (i.e., $K_{\theta}\propto a^{-2D}$) would, if it exists, last for only a short duration. Note also that, even though $K_{\theta}$ is the largest component of the energy density during much of the earlier evolution shown in Fig. 26 during which vortices are present (spanning around a decade of growth in $a$), the scaling $K_{\theta}\propto a^{-8/3}$ (i.e., $a^{-4D/3}$ for $D=2$) is not a period of kination-like scaling (i.e., $a^{-2D}$). [^38] In fact, in $D=3$, $a^{-4D/3}$ would be $a^{-4}$, the same scaling as radiation.
@@ 6 Phenomenological Implications: Periodic Gravitational-Wave Bursts
The amplitude of the observable gravitational-wave bursts at the peak frequency $f^{i}_{\Omega,{\rm obs}}$ depends also on the amplitude at emission, as well as redshifting effects. The gravitational-wave burst, once emitted, redshifts like radiation $h^{i}\propto(a_{i}/a)$. The peak amplitude of the individual burst, just like its frequency $f^{i}_{\Omega,{\rm emit}}$, can depend on the scale factor at emission $a^{i}$. Such a dependence stems from how the various (energy) densities and energy scales depend on $a$. As shown in Fig. 25, the total energy density in the system, as well as the gradient energy density in the domain boundary regions, both redshift as $a^{-4}$. On the other hand, the string number density redshifts as [^39] This statement follows from the peak values of $\xi$ shown in the center panel of Fig. 25 being approximately constant with $a$; because $\xi$ is a ratio of a comoving length to a comoving volume, it follows that the physical number density redshifts as $\sim\text{const.}\times a/a^{3}\propto a^{-2}$. $a^{-2}$ (ignoring the burst oscillations). This suggests different relations between the peak amplitudes of the forest of peaks. Whereas a dedicated numerical study will be needed to understand which sets the relative size of the peak amplitude, some qualitative understanding can be obtained. Let us focus on the case where the peak frequencies are related by (51) and assume that the spectrum of each burst $\Omega_{i}(f)$ scales as $a_{i}^{-Q}$ at the peak frequency. In this case, the total signal $\Omega_{{\rm obs},i}(f)$ will scale as $a_{i}^{4-Q}\propto(f^{i}_{\Omega,{\rm obs}})^{4-Q}$. Correspondingly, in the region where the individual peaks significantly overlap, the different gravitational-wave bursts sum up to a power-law that scales at high frequencies as
@@
Further comment:
The preceding discussion is based on the dynamics of the vortex production and annihilation because these periodic dynamics are associated with the evolution of regions of high energy density in the field; these are expected to have associated GW production. However, it is unclear whether these dynamics dominate the GW spectrum, and/or whether there are other GW components that are sourced merely by the overall evolution of the inhomogeneous scalar field at late times, such that the overall spectrum might have both periodic-in-frequency and broad features. Recall also that even unambiguously delineating the energy density that is associated with “vortices” vs. that in the “background field” is a conceptual and operational challenge in the context of an excited field with nonlinear interactions; in turn, it may be difficult to unambiguously associate GW emission with “vortices” vs. “the rest of the field”. Ultimately, the GW spectrum arising from the full field evolution must be studied numerically; this is a task we defer to future work.
@@ 7 Conclusion and Remarks
The results presented in this paper may also be useful for understanding preheating after inflation Dolgov:1989us ; Traschen:1990sw ; Kofman:1994rk ; Kofman:1997yn ; Tkachev:1998dc . [^40] We thank Sergey Sibiryakov for discussions on this point and pointing out Tkachev:1998dc , in which string formation was first observed in the study of preheating after inflation. Separately, an interesting question for future work is whether the dynamics we discovered in this paper survive if the $U(1)$ is weakly gauged, or if the $U(1)$ is extended to a larger symmetry group (e.g., the case of the Standard Model electroweak sector) Dufaux:2010cf ; Rajantie:2000nj ; Amin:2014eta . Would defects form periodically if the Standard Model Higgs were displaced in the radial direction, and would those dynamics remain synchronized across Hubble patches?
@@ Acknowledgments
We thank Will East, Sung-Sik Lee, Sergey Sibiryakov, and Zach Weiner for helpful discussions, and Anson Hook for comments on the draft. This work used anvil at Purdue University through allocation PHY240191 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program access , which is supported by National Science Foundation (NSF) Grants Nos. OAC-2138259, -2138286, -2138307, -2137603, and -2138296. J.H. would like to thank KITP for hospitality during the completion of this work. This research was supported in part by NSF Grant No. PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). M.A.F. thanks the Aspen Center for Physics for hospitality during early stages of this work, supported by NSF Grant No. PHY-2210452. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities.
@@ Appendix A Concise Review of Vortices in Condensed Matter Systems
$S=-\sum_{j}n_{s}m_{*}K_{j}\int X_{j}dY_{j}+\sum_{i\neq j}\frac{n_{s}m_{*}}{2\pi}K_{i}K_{j}\int\ln\left|\vec{R}_{i}-\vec{R}_{j}\right|dt\>,$ (59)
first discussed in 1985PhRvL..55.2887H ; 1993PhyA..200…42T , is an effective description of vortex interactions in the Ginzburg–Landau model in 2D (classical XY model). Here, $\vec{R}_{i}=\left\{X_{i},Y_{i}\right\}$ is the location of the vortex, while the circulation $K=\pm 2\pi/m_{*}$ for vortices (++) and anti-vortices (-). It was noted in 1993PhyA..200…42T that the first term in the action is a Berry phase and that this action resembles the action describing the motion of a charged particle (in the XY-plane) in an external magnetic field pointing in the Z-direction (see Tong:2016kpv for a detailed discussion about the connection with the quantum Hall effect). The resulting equation of motion of a vortex $i$ in the background of all the other vortices can be found to be
@@
Particle–vortex duality sheds light on this problem Peskin:1977kp ; Dasgupta:1981zz ; Karch:2016sxi ; Senthil:2018cru ; beekman2011electrodynamics . In two dimensions, particle–vortex duality can be established between the Abelian Higgs model and the XY model. Specifically, the Coulomb phase of the Abelian Higgs model is dual to the Higgsed phase of the XY model. The charged Higgs in the Abelian Higgs model is dual to the vortex in the XY model, and the gauge field is dual to the velocity field through Senthil:2018cru (see footnote 21 for the definition of the Levi-Civita tensor in 2+1-dimensions)
@@ Appendix B Simulation Details
respectively. The gradients of the former, just like $\theta$, are ill-defined at vortex locations, leading to numerical artifacts also in the gradient energy density $g_{\theta}$. The equations are discretized using fourth-order accurate finite-difference stencils in conjunction with fourth-order accurate Runge-Kutta time stepping. The grid spacing is $\Delta x$ and related to the timestep size $\Delta t$ by the Courant factor: $\Delta t=\Delta x/2$. Periodic boundary conditions are imposed in both two- and three-dimensional settings, where the box with volume $(2L)^{D}$ has linear size $L=10\mu^{-1}$ and $L=5\mu^{-1}$ for all two-dimensional simulations on a Minkowski and expanding FLRW background spacetime, respectively. Kreiss–Oliger numerical dissipation is used for numerical stability. This implies that modes of wavelength comparable to the grid spacing, $\lambda\sim\mathcal{O}(1)\Delta x$, are damped efficiently and removed from the evolution. This technique also automatically damps out unresolved (i.e., sub-grid/high-momentum) modes in the initial conditions after some transient; we find these transients to be irrelevant to our conclusions (see also comments at footnote 15). The scale-invariant Gaussian random field, entering the initial conditions described in detail in Sec. 2.4, is constructed using the Box–Muller approach. Lastly, vortices are located using the algorithm introduced in Fleury:2015aca . In two spatial dimensions ($D=2$), this unambiguously identifies the vortices piercing a cell. In three spatial dimensions ($D=3$), we search for strings through only a single face of a given cell (say, the face orthogonal to the +x+x-direction) to find the following lower bound on the string length density $\xi$: this density $\xi=\Delta x/V_{b}(N_{S}/N_{T})$ is obtained by comparing the total number of cells, $N_{T}$, to the total number of cells with a face pierced by a string, $N_{S}$, where the total simulation volume is $V_{b}$. This generally underestimates the length density, since the length of a string’s segment associated with a cell is at least [^42] Because the string cores are well-resolved, the string cannot be highly curved within a single cell. While there may be a small, $\mathcal{O}(1)$ inaccuracy in these per-cell length bounds owing to string curvature, this is likely a negligible effect. $\Delta x$. In principle, the length of this segment could be as long as the simulation box’s linear size $2L$ if the string is aligned with the faces. However, in practice, the string network lacks order on scales comparable to the box size. Therefore, the density $\xi$ is likely at most a factor of a few below the true string length density. In those cases with complete string annihilation, the absence of strings piercing all six faces of each cell in the simulation domain was checked.
@@
Subtleties arise when determining the critical radial displacements $\rho_{c}$ as a function of $\mathcal{J}_{0}$ shown in Fig. 12. Not only could the threshold depend on resolution as outlined in the above discussion, but due to the small number density of vortices for $\rho_{i}\gtrsim\rho_{c}$, the critical threshold $\rho_{c}$ may depend on the noise realization $\mathcal{N}(\bm{x})$. Precisely identifying this threshold requires considering an average over a series of noise realizations, as well as varying $\sigma_{\rho_{i}}$. In this work, we simply obtain approximate thresholds as follows. At fixed $\mathcal{J}_{0}$, we perform a suite of simulations spanning a range of initial displacements $\rho_{i}$ from just below to just above the threshold $\rho_{c}$. For each such displacement, we consider only a single noise realization and fix the variance to $\sigma_{\rho_{i}}=10^{-6}\rho_{i}$. We then estimate the threshold as the mean $\rho_{c}\approx(\rho_{i}^{<}+\rho_{i}^{>})/2$, between the initial displacement $\rho_{i}=\rho_{i}^{<}$, which marginally produced vortex pairs, and $\rho_{i}=\rho^{>}_{i}$, which lead to no vortices throughout its evolution. The errorbars shown in Fig. 12 are then simply $\pm 2(\rho_{i}^{>}-\rho_{i}^{<})$. Importantly, we do not strictly rule out that vortices may form for $\rho_{\text{min}}<\rho_{i}<\rho_{i}^{<}$, due to the subtleties listed above.
@@
For the comparison of the growth rates of the various spatial Fourier modes, shown in the right panel of Fig. 7 for $D=2$, we proceeded as follows. At each time step throughout the evolution, we compute the spatial Fourier transform $\tilde{\rho}(t,\bm{k})$ of $\rho(t,\bm{x})-\langle\rho(t,\bm{x})\rangle$ and transform to cylindrical coordinates $\bm{k}=(k_{x},k_{y})\rightarrow(k_{r},k_{\varphi})$. Since the spatial dependence of $\rho(t,\bm{x})-\langle\rho(t,\bm{x})\rangle$ is initialized to be a scale-invariant Gaussian random field, $\tilde{\rho}(t,\bm{k})$ exhibits a high degree of spherical symmetry in momentum space. We define the angular-averaged Fourier transform $\tilde{\rho}_{\varphi}(t,k_{r})=(2\pi)^{-1}\int_{0}^{2\pi}dk_{\varphi}\,k_{r}\tilde{\rho}(t,k_{r},k_{\varphi})$. Finally, for a given $k_{r}$-mode, we fit an exponential $\sim\exp(\Gamma t)$ to $\tilde{\rho}_{\varphi}(t,k_{r})$ to determine $\Gamma(k_{r})$. Notice, the value $\Gamma(k_{r})$ depends on the fitting time window as follows. At early times, $\tilde{\rho}_{\varphi}(t\mu\ll 1,k_{r})$ is constant in $k_{r}$ (up to fluctuations varying between noise realizations), sourcing all unstable modes at roughly the same amplitude. After a few e-folding times of the most unstable mode, the amplitudes of these modes differ by orders of magnitude. This implies that the numerical discretization error associated with the most unstable mode mixes into those unstable modes with smaller $\Gamma(k_{r})$. Therefore, at late times (before the modes reach the nonlinear regime) most $k_{r}$-modes grow with similar growth rates dictated by the most unstable mode. This is a numerical artifact and can be remedied to a degree by increasing $L$ and spatial resolution. To obtain $\Gamma(k_{r})$ for the case presented in the right panel of Fig. 7, we restrict to a fitting window early on in the exponential growth of $\tilde{\rho}_{\varphi}(t,k_{r})$, where the growth rate of $k_{r}$-modes with smaller growth rates are not significantly affected by the truncation error of the most unstable mode. The errorbars shown in Fig. 7 are determined taking into account different fitting windows, a simulation with $\times 2$ the resolution, and a simulation with $\times 4$ the box volume. Notice, this uncertainty is larger for those $k_{r}$-modes with smaller $\Gamma(k_{r})$. Finally, extracting $\Gamma(k_{r})$ from $\theta(t,\bm{x})-\langle\theta(t,\bm{x})\rangle$ yields results consistent with those presented in the right panel of Fig. 7.
@@
To test our cosmological implementation, we solve for the evolution of the zero-mode $\rho_{0}$ governed by (10), augmented by the appropriate Hubble friction in two spatial dimensions, as shown in (44). This simpler set of ordinary differential equations can be readily solved using standard numerical integration techniques. The solution can be compared with the early-time (i.e., before unstable modes acquire a significant amplitude) solution of a cosmological simulation solving the full complex scalar field equations on a radiation-dominated FLRW background spacetime. For the purposes of this comparison, we set $\sigma_{\rho_{i}}=10^{-14}\rho_{i}$ to ensure the unstable modes are sourced at sufficiently small amplitudes. In Fig. 28, we compare these two solutions. Clearly, at early times these two solutions are in excellent agreement. Furthermore, with increasing resolution of our fully nonlinear numerical methods, the difference converges at the expected fourth order. To convince ourselves that the vortex annihilation and transition to the radiation state at late times during the cosmological evolution is not due to insufficient resolution, we show the relevant energy densities in the left panel of Fig. 28. Evidently, the densities are converged in the relevant regime, and the transition to radiation is independent of resolution. The curves only begin deviating from each other once $(a\mu)^{-1}$ approaches the grid scale. In the bottom panel of Fig. 28, we show the change of the vortex density with resolution. The periodicity is a consistent feature across resolutions, while the overall density may vary from burst to burst between resolutions (for all the reasons outlined above, including—but not limited to—necessarily different random realizations of the initial conditions [i.e., phases and amplitudes] across the different-resolution simulations). Lastly, the bursts of vortex production in different-resolution simulations experience a mild dephasing between early and lates times. This originates both from an initial phase-offset at the time of the first burst, as well as numerical drift accumulated throughout the simulations. However, the total dephasing of the bursts accumulates to $\sim\mathcal{O}(1)$ radians across $\sim\mathcal{O}(100)$ bursts.